# Counting normals to closed curves in $\mathbb{R}^{3}$

### v1

@@ Abstract.
(1) For every generic closed smooth curve in $\mathbb{R}^{3}$, there is a point with at least $6$ emanating normals to the curve. If the curve is knotted, there is a point with at least $8$ emanating normals.
@@ 1. Preliminaries and basic definitions
If the curve is knotted, there is a point with at least $8$ normals.
@@
Throughout this section we assume that $C$ is a generic closed smooth curve embedded in $\mathbb{R}^{3}$. From now on, we call local maxima and local minima simply maxima and minima, respectively, for brevity.
@@
For $x\in C$ let us fix the following notation: $h(x)$ is the osculating plane to $C$ at the point $x$, the point $F(x)\in h(x)$ is the center of the osculating circle. A generic curve has an everywhere non-zero curvature, so we always assume that $F(x)$ is not at the infinity.
$L(x)$ is the line passing through $F(x)$ and orthogonal to $h(x)$. That is, $L(x)$ goes in the direction of the binormal $b(x)=t(x)\times n(x)$. The focal surface $\mathcal{F}(C)$ is the union of all these lines. We refer the reader to [2] and [10] for more details on spatial curves.
We recall that $\mathcal{F}(C)$ is the bifurcation set of the function $SQD_{y}$. That is, $\mathcal{F}(C)$ cuts the space into chambers. In each of the chambers the number of normals $N(y)$ persists, and if $y$ crosses a sheet of $\mathcal{F}(C)$ transversally, either two normals are born, or two normals die.
Proposition 1.
The focal surface $\mathcal{F}(C)$ of any generic closed smooth curve in $\mathbb{R}^{3}$ has a transverse double self-intersection.
Proof.
The focal surface $\mathcal{F}(C)$ is equal to the tangent developable of the evolute $\mathcal{E}(C)$, that is, $\mathcal{F}(C)$ is the union of all tangent lines to $\mathcal{E}(C)$ [2]. The evolute is generically a closed curve, possibly with cusps and points at infinity.
  1. (1)
    Assume that the torsion of $C$ never vanishes. Then $\mathcal{E}(C)$ has no points at infinity. Consider the following two cases:
    (a) $\mathcal{E}(C)$ has no cusps. We have to prove that that $\mathcal{E}(C)$ has two intersecting tangent lines. Take a support plane to the curve $\mathcal{E}$. It is a plane that is tangential at some point $x\in\mathcal{E}$ and intersects $C$ at no other point. Start rotating the plane around the tangent line to $\mathcal{E}(C)$ at the point $x$. Sooner or later the plane hits $\mathcal{E}(C)$ at some other point $x^{\prime}$. Both tangent lines (to $x$ and to $x^{\prime}$) lie in the plane and therefore generically intersect.
    (b) Assume that $\mathcal{E}(C)$ has a cusp. The cusp corresponds to a swallow tail of $\mathcal{F}(C)$, which provides a curve of transverse self-intersection.
  2. (2)
    Assume that the torsion of $C$ changes the sign at some point $x_{0}\in C$. We shall prove that there is a pair of intersecting lines $L(x)$ and $L(x^{\prime})$ for some $x$ and $x^{\prime}$ lying near $x_{0}$ (in fact, we show that there is a one-parametric family of such pairs).
    An ordered pair of oriented lines in $\mathbb{R}^{3}$ is skew if they neither intersect nor are parallel. Each skew pair of lines has a naturally defined orientation: we have an ordered triple of vectors, the defining vectors of the lines, and the vector associated to their common perpendicular which is directed from the first line to the second. Then, by definition, the orientation of the pair of lines is the orientation of the triple of vectors. The orientation clearly persists under continuous deformations of the pair provided that the lines remain skew during the deformation.
    For $x\in C$ let us orient the line $L(x)$ the direction of the binormal. Pick two different points $x,x^{\prime}\in C$ close to $x_{0}$ and lying on one side to $x_{0}$. The orientation of the associated pair $L(x),L(x^{\prime})$ depends on the sign of the torsion. Let the pair $x,x^{\prime}$ move along the curve to some target pair of points on the other side with respect to $x_{0}$. Since the orientation of the pair $L(x),L(x^{\prime})$ changes, somewhere in between the lines intersect. This corresponds to self-crossing of $\mathcal{F}(C)$.
Now we are ready to prove Theorem 1:
Proof.
A neighborhood of a point of transverse self-intersection of $\mathcal{F}(C)$ is divided into four chambers by two sheets of $\mathcal{F}(C)$. Crossing of each sheet changes $N(y)$ by $2$. Each sheet can be given a co-orientation in the direction of the increasing number of normals. It follows, that in these four chambers we have $N$, $N+2$ and $N+4$ emanating normals for some $N$. The value of $N$ is at least $2$, whereas in the knotted case by Lemma 1 it is at least $4$ which implies the desired. ∎
We present below a sketch of yet another proof of the first statement of Theorem 1, based on a different idea.
Proof.
@@
Remark. For closed PL-curves the number of normals is always finite and one has $N(y)\leq 2v$, where $v$ is the number of vertices.

### v2

@@ Abstract.
(1) For every generic closed smooth curve in $\mathbb{R}^{3}$, there is a point with at least $6$ emanating normals to the curve.
@@ 1. Preliminaries and basic definitions
There arises a natural question: Is it true that for every generic smooth knotted closed curve in $\mathbb{R}^{3}$, there exists a point $y$ with at least $8$ emanating normals?
@@
Throughout the paper we assume that $C$ is a generic curve embedded in $\mathbb{R}^{3}$. From now on, we call local maxima and local minima of $SQD_{y}$ maxima and minima, for brevity.
@@
For a smooth curve $C$ and a point $x\in C$ let us fix the following notation: $h(x)$ is the osculating plane, $x^{\perp}$ is the normal plane, the point $F(x)\in h(x)$ is the center of the osculating circle. A generic curve has an everywhere non-zero curvature, so we may assume that $F(x)$ is not at the infinity.
$L(x)$ is the line passing through $F(x)$ and orthogonal to $h(x)$. That is, $L(x)$ goes in the direction of the binormal $\textbf{b}(x)=\textbf{t}(x)\times\textbf{n}(x)$. The focal surface $\mathcal{F}(C)$ is the union of all these lines. We refer the reader to [2] and [9] for more details on spatial curves.
$\mathcal{F}(C)$ is the bifurcation set of the function $SQD_{y}$. That is, $\mathcal{F}(C)$ cuts the space into chambers. In each of the chambers the number of normals $N(y)$ persists, and if $y$ crosses a sheet of $\mathcal{F}(C)$ transversally, either two normals are born, or two normals die.
Here is the proof of Theorem 1:
@@
Remark. Assume that $C$ is knotted, and the focal surface $\mathcal{F}(C)$ has a transverse self-intersection. Then there is a point with at least $8$ emanating normals.
Indeed, a neighborhood of a point of transverse self-intersection of $\mathcal{F}(C)$ is divided into four chambers by two sheets of $\mathcal{F}(C)$. Crossing of each sheet changes $N(y)$ by $2$. Each sheet can be given a co-orientation in the direction of the increasing number of normals. It follows, that in these four chambers we have $N$, $N+2$ and $N+4$ emanating normals for some $N$. The value of $N$ is at least $4$ which implies the desired.
@@
Remark. For closed piecewise linear curves the number of normals is always finite and one has $N(y)\leq 2v$, where $v$ is the number of vertices.